what is the answer to 4/3u-6=7/3u+8

Answer by Alan3354(69443)

what is the answer to 4/3u-6=7/3u+8
--------------
Depends. Is it 4/(3u-6)? Or (4/3u) - 6 = ??
Parentheses are free, use some.